A steam engine operates between 500oC and 300oC. What is the maximum possible efficiency?
Answer #1

efficiency of a carnot engine is,

e=1-TC/TH

TH=5000C=500+273=773 K

TC=3000C=300+273=573 K

e=1-573 K/773 K

=0.258

efficiency in percentage is 25.8%

approximately 26%